Blood serum is typically about 0.0044 M in K+ (aq). Calculate the mass of the potassium ions in 4,459. mL of blood serum. The molar mass of potassium is 39.10 g/mol.
1
Expert's answer
2019-06-26T04:03:20-0400

Firstly, we should count out the amount of K in the solution. We have the molarity of the solution (c), which is 0.0044 M or 0.0044 mol/l. Now we will convert the volume of the solution (v) from ml to l.

4.459 ml = 4.459 * 10-3 l;

n = c * v = 0.0044 * 4.459 * 10-3 = 4.4 * 10-3 * 4.459 * 10-3 = 1.96 * 10-7mol;

n = m/Mr, that is why:

m = n * Mr = 1.96 * 10-7 * 39.1 g = 7.66 * 10-8 g = 76 ng;